PREX/CREX Apr 30, 2019 10:00am

From PREX Wiki
Jump to: navigation, search

Back to Main Page >> PREX/CREX_Teleconferences

previous meeting << >> following meeting

Logistic information

 BlueJeans calling instructions:
 Toll-Free Number (U.S.&  Canada):  888-240-2560
 International toll number:         408-740-7256
 Bluejeans CODE:                  628 981 532
 Bluejeans link: https://bluejeans.com/628981532
 Offline room:   CEBAF Center F324-325

Agenda

  1. Installation progress: Ciprian docDB363
    • SR: is 0.1 deg important. KP: it’s not nothing but it’s close to be within specs. I estimate this difference fewer than 1mm at the pinch point. We need to see how it fits. If they pull out several cm downstream of the flanges we could get clipping before the Q1 collimator.
    • SR: got an email from Jack and he is predicting he will be ready by next Tue.
    • NL: do we have an assigned location for constructing the GEMs? SR: we should ask for space.
  2. Compton polarimeter
    • DG: The 10W fiber amp died and has been repaired and should be shipped back soon. We are using the 5W fiber amp to look at polarization effects. We let out vacuum and are doing further investigation.
  3. Moller polarimeter
    • SP: We have the Cu dummy target and adding some insurance for the rotation. A screw stop has been added to fix the target if we don’t need to it. Eric is working on the PREX tune and compare with the G3. Further looking at the data analysis. We have a new PMT for the scintillatior; it is being now tested. Preparing run plan for PREX.
  4. BPM studies hapLog3738
    • CP: Att1 spring data at 240Hz. There are jumps that occur every second. The thought is that is we keep the switch clock constant we won’t see any jumps.
    • CP: BPM12 is always in FastFeedBack and we read those out. We could move to this setting for all of our BPMs. However several other signals are changed, and OPS “sees” jumps (because the slow feedback works not as well).
  5. DAQ move to adaq1
    • BM: Sat (may the 4th) is the planned move of the daq. It may mean that some things won’t work -- will try to fix all scripts. Data will be written to the local disk (adaq1). KP: we just need to update the env variable. CG: i’ll make sure that the links are ok
  6. LQ1 viton seals:
    • CG: Adam has an updated configuration and we hope to finalize this issue this week so that we can communicate with Jessie if we want additional shielding.
  7. Injector studies:
    • CP: we will need to put in batteries and look if there is any pickup in the injector.
  8. Counting DAQ update:
    • BM: Chandan and I setup some triggers and we have a basic DAQ working. We have a basic pod going. We hope to look at the pulser later this week. We are looking at the Left arm and hope to repeat this with the Right.
  9. Discussion for next week:
    • CG: if you finalize anything please document everything and present the summary at this meeting

Present

In F324

Paul K, Sanghwa P, Caryn P, Ciprian G, Adam Z, Tao Y, Devi A

Online

Ye T, Juan Carlos C, Eric K, Kent P, Dave G, Victoria O, Sakib R, Jay B, David A, Seamus R, Rich H, Kent P, Amali P, Chandan G, Guido U, Bob M, Alexandre C, Dustin M, KK, Ryan R, Cameron C, Nilanga L, Paul S, Wouter D, Siyu J,

Cannot make it